What is [H3O+] for normal aspirin?

What is [HO-] in distilled water?

What is [H3O+] for shampoo?

What is the pOH of tap water?

What is the pOH of normal aspirin?
Answer #1

Regular aspirin will dissolve in water, which has a pH of approximately 7.0. This pH is found in the mouth.

Regular aspirin or buffered aspirin will dissolve in acid. An acidic pH is found in the stomach (pH < 4.0).

Enteric aspirin will dissolve in a base. A basic pH is found in the intestines (pH > 8.0).

As aspirin is a weak acid, [H3O+ ] must be calculated to find pHHAsp(aq) + H2O(1) $ H20(aq) + Asp (aq)

So ,pH and [H3O+] are related as [HO]= 10-pH . pH of asprin depends upn the cncentration of its solution at a particular temperature.

  Because,pH + pOH = 14,therefore you can find pOH of normal aspirin after Calculating the pH value.

For Distilled water and tap water, the pH value can vary from 0 to 14. As we know,Solutions with a pH between 0 and 7 are acidic, while those with a pH between 7 and 14 are basic. Pure distilled water should be neutral with a pH of 7, but because it absorbs carbon dioxide from the atmosphere, it's actually slightly acidic with a pH of 5.8.

To a very small but very important extent, water is a weak acid that ionizes to hydrogen and hydroxide ions.

H2O + H+ + OH

In pure water, the hydrogen ion concentration, [H+], equals the hydroxide ion concentration, [OH-]. These concentrations can be calculated from the equation for the ionization of water.Let x equal the hydrogen ion concentration, [H+]. Then x also equals the hydroxide ion concentration, [OH-].

[H+][OH-] = (x)(x) = 1 X 10-14
x2 = 1 X 10-14
x = 1 X 10-7
[H+] = [OH-] = 1 X 10-7 M

Hydroxide ion concentration in an acid solution the following relationships exist whenever water is present,Suppose the hydrogen ion concentration is 0.10 M,then

[H+][OH-] = 1 X 10-14 and pH + pOH = 14

[OH-] = 1.0 X 10-14 /0.10 = 1 X 10-13 M

Suppose the pH of an acidic solution is 1.0.Then,

1.0 + pOH = 14      pOH = 13

If ,A shampoo has a pH of 8.56,then

[H3O+] = 10^-8.56 = 2.7 × 10^-9 M